OpenStudy (anonymous):

3.Jesse rents a car for an initial $75 for the first 50 miles. However, she must pay $2 for every mile she drives. The following week, Alex rents the same car and pays $80 for the first 50 miles and $1.50 for each additional mile. If they each paid the same amount and drove the same number of miles, how far did each travel?

OpenStudy (anonymous):

75 + 2x = 80 + 1.5x

OpenStudy (anonymous):

is that the formula, like to set up the problem or is that the answer?

OpenStudy (anonymous):

x is the number of additional miles each person drove\[75+2x=80+1.5x\]\[2x-1.5x=5\]\[x(2-1.5)=5\]\[x=\frac{5}{(2-1.5)}\]\[x=\frac{5}{0.5}\]\[x=10\]So the total number of miles each drove is 50 + x, or 50 + 10, or 60 miles.